OpenStudy (anonymous):

Easy, but I forget how to do it: 2 x -- = ----- 3 12000

OpenStudy (anonymous):

x=8000

OpenStudy (anonymous):

How do you do it?

OpenStudy (anonymous):

2/3=x/12000 you gotta times the 12k or switch to the other side 2x12000=24k 24000/3=8000

OpenStudy (anonymous):

Ah right, thanks